OpenStudy (anonymous):

Find the distance between P and Q. If neccessary round to the nearest hundredth. 1. P(-3, -2) and Q(3, 6) 2. P(-2, 7) and Q(-6, 4) 3. P(-4, -13) and Q(16, 8) 4. P(-2, 5) and Q(8, 7)

OpenStudy (anonymous):

iamgaygaygaygay: Damn Bastards!Damn Bastards!Damn Bastards!Damn Bastards!Damn Bastards!Damn Bastards!Damn Bastards!Damn Bastards!Damn Bastards!Damn Bastards!

OpenStudy (hoblos):

distance between two points (x1,y1) and (x2,y2) is \[D = \sqrt{(x_2-x_1)^2+(y_2-y_1)^2} \] can you apply it and tell me what you get

OpenStudy (anonymous):

iamgaygaygaygay: Damn Bastards!Damn Bastards!Damn Bastards!Damn Bastards!Damn Bastards!Damn Bastards!Damn Bastards!Damn Bastards!Damn Bastards!Damn Bastards!

OpenStudy (anonymous):

number one would be 4 correct?

OpenStudy (hoblos):

not really let's do it step by step P(-3, -2) and Q(3, 6) x2-x1 = 3-(-3) = ?

OpenStudy (anonymous):

why did you put the -3 behind? why dont you put it first?

OpenStudy (hoblos):

you can do that so it will be -3-3 = ?

OpenStudy (anonymous):

6 correct?

OpenStudy (anonymous):

so once i have added it all up the final answer is 91 right?

OpenStudy (hoblos):

if you use 3-(-3) it would be 6 if you use -3-3 it would be -6 but for both cases you will get the same answer for the distance since you will square it now

OpenStudy (anonymous):

so it would be 9.5........

OpenStudy (hoblos):

umm let's continue it step by step to get the correct answer 6^2 = ?

OpenStudy (anonymous):

36

OpenStudy (hoblos):

correct, now for the y y2-y1 = 6-(-2) =?

OpenStudy (anonymous):

=8 8^2= 64

OpenStudy (hoblos):

correct 36+64=?

OpenStudy (anonymous):

100 square root of 100=10

OpenStudy (hoblos):

correct answer

OpenStudy (anonymous):

so i just do the same for them all?

OpenStudy (anonymous):

thank you so much!

OpenStudy (hoblos):

yeah do the same for the others you can tell me what you get to check the answers for you

OpenStudy (anonymous):

so the next one would be 5 correct?

OpenStudy (hoblos):

correct

OpenStudy (anonymous):

and the one after that would be 29?

OpenStudy (hoblos):

correct too

OpenStudy (anonymous):

and the last one is 10.19?

OpenStudy (hoblos):

do you have choices for the answer or you have to write it ?

OpenStudy (anonymous):

write it down

OpenStudy (hoblos):

ok so if we apply the formula we get radical(104) = 10.19803... so if you want to round it to the hundredth it would be 10.2 and if you want to round it to the thousandth it would be 10.198

OpenStudy (anonymous):

okay thank you!

OpenStudy (hoblos):

glad to help! and good work :)